# Real-time Event Recognition of Long-distance Distributed Vibration Sensing with Knowledge Distillation and Hardware Acceleration

### v1

@@ Abstract
Distributed optical fiber vibration sensing (DVS) technology based on phase-sensitive optical time-domain reflectometry is widely used for safety monitoring and intrusion event surveillance in wide-ranging fields. Existing methods rely on deep learning models for event recognition but struggle with real-time processing of large data volumes in long-distance applications. To address these challenges, we use a four-layer convolutional neural network (CNN). The application of knowledge distillation with ResNet as the teacher model improves the generalization ability of the four-layer CNN, increasing the accuracy from 83.41% to 95.39% on data from untrained environments. The model is implemented on a field programmable gate array (FPGA) using a novel design that replaces multiplication with binary shift operations and quantizes model weights accordingly, allowing for high parallelism and low latency. An inference time of 0.083 ms is achieved for a spatial-temporal sample with a 12.5 m fiber length and 0.256 s time frame. This implies the system can process signals over a fiber length of approximately 38.55 km in real time, which is more than twice the capability of a GPU of Nvidia GTX 4090. The proposed method greatly improves the efficiency of vibration pattern recognition, thus promoting the application of DVS as smart sensing system in various areas. The data and code is available at https://github.com/HUST-IOF/Efficient-DVS.
@@ I Introduction
Distributed optical fiber vibration sensing (DVS) technology is based on phase-sensitive optical time-domain reflectometry (phi-OTDR) system, and uses optical fibers as the sensing element. It is capable of providing dense and continuous wide-range measurement of vibration from surrounding environment and recognize various intrusion events based the sampled data. This technology has been applied as smart sensing system in various areas in recent years, including railway safety monitor[1, 2], perimeter security[3, 4], pipeline monitor[5, 6], and smart city[7]. As shown in Fig. 1, the phi-OTDR system captures the backscattering Rayleigh light across the fiber. The collected Rayleigh light trace, whose phase can be affected by external vibration, provides valuable information about the vibration events.
@@
Signal processing algorithms are able to extract information from vibration traces. However, the recognition accuracy of vibration events is often limited by the variability of signals caused by complex and changing environments, as well as the diverse types of intrusion events encountered in practical applications. To improve the recognition performance, conventional methods typically involve hand-crafted algorithms for data preparation and feature extraction, followed by the use of classic machine learning algorithms for recognition[8, 9, 10, 11, 12, 13]. However, this kind of approaches can be time-consuming in both development and deployment. Designing feature extraction algorithms requires researchers to possess knowledge of signal processing algorithms and the specific application context. As a result, the resulting algorithms can be complex, time-consuming, and may not capture the most important features accurately.
To address these challenges, deep learning models (DLMs) have been introduced to the field. Initially, DLMs were introduced for pattern recognition based on extracted features [9]. With the development of more sophisticated and computationally powerful DLMs, they are capable of automatically extracting important features from noisy data and simultaneously handling feature extraction and pattern recognition tasks. The introduction of complex DLMs brings about improved performance and reduces the design workload. Consequently, they have received more attention in recent years. Table I shows several typical methods of this type of application, where $GAF$ represents gramian angular field, $S-T$ denotes space-time, and $F-T$ is frequency-time.
@@
However, the utilization of high-capacity DLMs has resulted in increased computation resource requirements. Moreover, when using the DVS system to monitor optical fibers, it is important to note that the volume of captured data directly relates to the length of the fiber. Due to the limited available computation resources in real-world applications, achieving real-time recognition is only possible over short-distance fibers [17]. To improve the computational efficiency of the DVS system, there are typically three solutions. The first is data selection or down-sampling, which involves using prior knowledge and data analysis to identify and retain valuable signal sections or sample points[2, 7]. This approach aims to reduce the input data volume through hand-crafted algorithms, though it may result in false negatives. The second solution is model compression, which seeks to lower computational complexity and resource demands. However, existing solutions use models with a parameter number no lower than millions to ensure performance and generalizability[7]. So, there is still potential for further improvement to create smaller and more efficient models. The third solution is the use of high-performance hardware, such as graphics processing units (GPUs) and cloud computing. While this approach can significantly increase computational power, it is often expensive and demands high power consumption, as well as a reliable communication system to handle the large volume of sample data.
@@
This paper proposes to improve the signal processing capability of the DVS system through knowledge distillation and hardware acceleration. We propose to use a shallow convolutional neural network (CNN), instead of DLMs, to reduce model complexity and computation resource requirements. Moreover, we identify the generalizability limitation of low-capacity models and propose to use the knowledge distillation technique as a solution. To further improve the throughput of the data processing system, we design an efficient hardware acceleration scheme based on field programmable gate array (FPGA). Additionally, we have integrated an Analog-to-Digital Converter (ADC) with the FPGA to minimize latency during data transfer, as depicted in Fig. 1. The proposed scheme utilizes FPGA’s flexibility to achieve full optimization of the model. It involves replacing multiplication operations in the model inference with shift operations, as the later ones can be implemented without digital signal processing (DSP) resources. This design is only limited by the logic resources, instead of DSP resources which is more limited[18], on the chip, allows for higher parallelism and thus faster calculation. This architecture is referred as the shift-add structure in the rest of the paper. Additionally, we propose to quantize the model weights into integers that indicate the number of binary shift operations. The quantization technique is referred as shift parameter quantization in the rest of the paper. Through this approach, we can achieve real-time processing of a large amount of captured data from the DVS system on a mainstream industrial FPGA.
@@
  • Shallow CNN models, coupled with knowledge distillation, are used to develop a lightweight model that offers high performance and generalizability. This approach effectively addresses the throughput limitation brought by the high computational complexity of existing algorithms.
  • A novel hardware acceleration scheme targeted for shallow CNN is developed, leveraging shift-add structures and shift parameter quantization to achieve higher efficiency.
@@ II-A Knowledge Distillation based on logits
where $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]$ is the input vector, which is the output of the network in this case. The knowledge is then transferred by guiding the student model to imitate the behavior of the teacher model. The process is to minimizing a loss function that measures the difference between the outputs of the teacher network and the student network, and also the difference between the outputs of the student network and the correct results, as shown in Fig. 2. The process can be seen as minimizing the difference between the outputs of the student network and the correct results. It can be achieved through minimizing a loss function that measures the difference between the outputs of the teacher network and the student network. The loss function usually consists of a standard classification loss, like cross-entropy, and a distillation loss that measures the difference between the softened probabilities generated by the teacher and student networks. Formally, the loss function $\mathcal{L}$ can be expressed as:
@@ II-B2 Activation Buffering
Obviously, both convolution layers and pooling layers only require the input feature map data within the selected window. Assume the input to be a spatial-temporal figure, where rows represent discrete times and columns represent positions. The data is streamed in row-major order. Then the activation buffer only needs to store $P$ rows from the input feature map during the computation process. Thus, the buffer mechanism is designed as shown in Fig. 4.
@@
When new data is inputted, the corresponding column shifts the existing data up by one unit, and the new data is filled into the bottom unit. This design allows the buffer to store just $P-1$ rows and $Q$ data from the last row to start the convolution operation. Thus, the design only requires $P$ rows of storage capacity. Moreover, reading new data and calculation can be performed in parallel to achieve lower latency. The convolution layer only needs to buffer $(P-1-S)*W+Q$ data for each channel instead of the complete input feature map to start operation, where $W$and $S$ denote the width of input feature map and the stride respectively. Thank you for the clarification. I apologize for the misunderstanding. In the case of a convolution layer, it only needs to buffer $(P-1-S)\times W+Q$ data for each channel, rather than the entire input feature map, in order to commence the operation. Here, $W$ represents the width of the input feature map, and $S$ represents the stride.
@@ II-B3 Shift-Add Structure
The main operation of convolutional neural networks is multiplication and accumulation (MAC). Digital signal processing (DSP) block is required to implement MAC units on FPGA. The limited number of DSP blocks restricts the number of MAC units that can be implemented on-chip. As a result, the throughput of the design is restricted. To avoid this restriction, it is proposed to use the shift and accumulation units that do not require DSP resources for implementation to replace the MAC units, as shown in Fig. 5. Digital values can be represented as a sum of powers of 2, as illustrated in Equation 5 with exponents denoted as $s_{0},s_{1},\ldots,s_{k}$. Multiplication with power of two is essentially performing shift operations on multiplier. This allows a MAC operations to be losslessly replaced by a series of shift and accumulation operations, as shown in Equation 6.
@@ III-A Performance evaluation
To evaluate and demonstrate the performance of the lightweight model, a 4-layer CNN with a structure outlined in Table III is utilized. The model consists of four layers and has a total of 30,771 parameters. In terms of computational cost, it requires 2,282,496 floating-point operations (FLOPs). Additionally, this 4-layer CNN model serves as the baseline for evaluating the improvements in both performance and generalizability achieved through knowledge distillation.
@@ IV Conclusion
This paper addresses the throughput limitation issue faced by the DVS systems based on DLMs, which prevents the implementation of real-time recognition over long-distance fibers. To overcome this challenge, we propose to use lightweight models as a replacement for DLMs to reduce the computation resource requirements of the system. We identify the problem of limited generalizability of the lightweight models and adopts knowledge distillation techniques to enhance the generalizability of the utilized lightweight model. Furthermore, we propose an FPGA-based hardware acceleration scheme for fast inference. It results in an impressive 2.14x shorter inference time compared to the state-of-the-art GPU implementation, and 4.52x compared to the CPU implementation, which is also the fastest recognition time reported so far in this field. Through our proposed method, we achieve real-time recognition over 38-km-long fibers. The implemented system allows for real-time recognition over long-distance fibers in practical applications, contributing to the advancement of DVS technology and its use as a smart sensing system in various areas.

### v2

@@ Abstract
Fiber-optic sensing, especially distributed optical fiber vibration (DVS) sensing, is gaining importance in internet of things (IoT) applications, such as industrial safety monitoring and intrusion detection. Despite their wide application, existing post-processing methods that rely on deep learning models for event recognition in DVS systems face challenges with real-time processing of large sample data volumes, particularly in long-distance applications. To address this issue, we propose to use a four-layer convolutional neural network (CNN) with ResNet as the teacher model for knowledge distillation. This results in a significant improvement in accuracy, from 83.41% to 95.39%, on data from previously untrained environments. Additionally, we propose a novel hardware design based on field-programmable gate arrays (FPGA) to further accelerate model inference. This design replaces multiplication with binary shift operations and quantizes model weights, enabling high parallelism and low latency. Our implementation achieves an inference time of 0.083 ms for a spatial-temporal sample covering a 12.5 m fiber length and 0.256 s time frame. This performance enables real-time signal processing over approximately 38.55 km of fiber, about 2.14×2.14$\times$2.14 × the capability of an Nvidia GTX 4090 GPU. The proposed method greatly enhances the efficiency of vibration pattern recognition, promoting the use of DVS as a smart IoT system. The data and code are available at https://github.com/HUST-IOF/Efficient-DVS.
@@ I Introduction
Distributed optical fiber vibration sensing (DVS) technology, based on phase-sensitive optical time-domain reflectometry (phi-OTDR), uses optical fibers as sensing elements to provide dense, continuous measurements of vibration and recognize various intrusion events. This technology has been applied as smart internet of things (IoT) systems in various areas in recent years, including industrial safety monitoring and intrusion detection. Recent applications of DVS technology include railway safety monitoring [1, 2], perimeter security [3, 4], pipeline monitoring [5, 6], and smart cities [7]. As shown in Fig. 1, the phi-OTDR system captures Rayleigh backscattering light across the fiber, with phase variations induced by external vibrations providing valuable event information.
@@
Despite its capabilities, DVS technology faces several challenges in IoT applications. The primary issue is the variability of signals due to complex and changing environments and diverse intrusion types, which limits event recognition accuracy. Traditional methods often rely on hand-crafted algorithms for data preparation and feature extraction, followed by classic machine learning for recognition[8, 9, 10, 11, 12, 13]. These approaches can be time-consuming and complex, requiring extensive expertise in signal processing. To solve these problems, deep learning models (DLMs) have been introduced to the field. Initially, DLMs were introduced for pattern recognition based on extracted features [9]. With the development of more sophisticated and computationally powerful DLMs, they are capable of automatically extracting important features from noisy data and simultaneously handling feature extraction and pattern recognition tasks. The introduction of complex DLMs brings about improved performance and reduces the design workload. Consequently, they have received more attention in recent years. Table I shows several typical methods of this type of application, where $GAF$ represents gramian angular field, $S-T$ denotes space-time, and $F-T$ is frequency-time.
@@
Another challenge in applying DVS technology in IoT applications is the large volume of captured data, which scales with the length of the fiber. Consequently, monitoring long-distance fibers requires processing vast amounts of data. Moreover, applications such as pipeline monitoring and earthquake detection demand real-time or near-real-time responses from smart IoT systems based on DVS technology to minimize potential property damage. In addition, the use of high-capacity DLMs significantly increases the demand for computational resources. In real-world scenarios, where resources are often limited, achieving real-time event recognition is typically feasible only for short-distance fiber monitoring [17]. To improve the computational efficiency of the DVS system, there are typically three solutions. The first is data selection or down-sampling, which involves using prior knowledge and data analysis to identify and retain valuable signal sections or sample points[2, 7]. While this reduces the input data volume, it can lead to false negatives due to its reliance on hand-crafted algorithms. The second solution is model compression, which seeks to lower computational complexity and resource demands. However, existing approaches often require models with at least millions of parameters to maintain performance and generalizability[7], leaving room for further optimization to develop smaller and more efficient models. The third solution is the use of high-performance hardware, such as graphics processing units (GPUs) and cloud computing. While this approach can significantly increase computational power, it is often expensive and demands high power consumption, as well as a reliable communication system to handle the large volume of sample data.
@@
This paper proposes enhancing the signal processing capabilities of DVS systems through the use of shallow convolutional neural network (CNN) coupled with knowledge distillation and hardware acceleration. We introduce a shallow CNN as a more efficient alternative to DLMs, reducing model complexity and computational resource requirements. we identify the generalizability limitation of low-capacity models and propose to use the knowledge distillation technique as a solution. To further improve the data processing throughput, we design an efficient hardware acceleration scheme based on a field programmable gate array (FPGA). An analog-to-digital Converter (ADC) is integrated with the FPGA to minimize latency during data transfer, as depicted in Fig. 1. The proposed scheme leverages FPGA’s flexibility to optimize model performance by replacing multiplication operations with shift operations, which are less resource-intensive. This shift-add structure, as we refer to it, allows for greater parallelism and faster computation, constrained only by the FPGA’s logic resources rather than its limited digital signal processing (DSP) resources[18] on the chip. Additionally, we propose a shift parameter quantization technique, where model weights are quantized into integers corresponding to the number of binary shift operations. This approach enables real-time processing of the substantial data volumes captured by the DVS system on a mainstream industrial FPGA.
@@
  • We propose to use shallow CNN models, coupled with knowledge distillation, to develop a lightweight model that offers high performance and generalizability. This approach effectively addresses the throughput limitation brought by the high computational complexity of existing algorithms.
  • We design a novel hardware acceleration scheme for shallow CNNs, utilizing shift-add structures and shift parameter quantization to significantly improve efficiency.
@@ II-A Knowledge Distillation based on logits
where $\mathbf{x}=[x_{1},x_{2},\ldots,x_{n}]$ is the input vector, which is the output of the network in this case. The knowledge is then transferred by guiding the student model to imitate the behavior of the teacher model. The process is to minimizing a loss function that measures the difference between the outputs of the teacher network and the student network, and also the difference between the outputs of the student network and the correct results, as shown in Fig. 2. The loss function usually consists of a standard classification loss, like cross-entropy, and a distillation loss that measures the difference between the softened probabilities generated by the teacher and student networks. Formally, the loss function $\mathcal{L}$ can be expressed as:
@@ II-B2 Activation Buffering
Obviously, both convolution layers and pooling layers only require the input feature map data within the selected window. Assume the input is a spatial-temporal figure, where rows represent discrete time points and columns represent spatial positions. The data is processed in row-major order. Then the activation buffer only needs to store $P$ rows from the input feature map during the computation process. Thus, the buffer mechanism is designed as shown in Fig. 4.
@@
When new data is inputted, the corresponding column shifts the existing data up by one unit, and the new data is filled into the bottom unit. This design allows the buffer to store just $P-1$ rows and $Q$ data from the last row to start the convolution operation. Thus, the design only requires $P$ rows of storage capacity. Moreover, reading new data and calculation can be performed in parallel to achieve lower latency. The convolution layer only needs to buffer $(P-1-S)*W+Q$ data for each channel instead of the complete input feature map to start operation, where $W$and $S$ denote the width of input feature map and the stride respectively.
@@ II-B3 Shift-Add Structure
The main operation of convolutional neural networks is multiplication and accumulation (MAC). DSP block is required to implement MAC units on FPGA. The limited number of DSP blocks restricts the number of MAC units that can be implemented on-chip. As a result, the throughput of the design is restricted. To avoid this restriction, it is proposed to use the shift and accumulation units that do not require DSP resources for implementation to replace the MAC units, as shown in Fig. 5. Digital values can be represented as a sum of powers of 2, as illustrated in Equation 5 with exponents denoted as $s_{0},s_{1},\ldots,s_{k}$. Multiplication with power of two is essentially performing shift operations on multiplier. This allows a MAC operations to be losslessly replaced by a series of shift and accumulation operations, as shown in Equation 6.
@@ III-A Performance evaluation
To evaluate and demonstrate the performance of the lightweight model, a 4-layer CNN with a structure outlined in Table III is utilized. The model consists of four layers and has a total of 30771 parameters. In terms of computational cost, it requires 2,282,496 floating-point operations (FLOPs). Additionally, this 4-layer CNN model serves as the baseline for evaluating the improvements in both performance and generalizability achieved through knowledge distillation.
@@ IV Conclusion
This paper addresses the throughput limitations of DVS systems, which hinder the implementation of real-time recognition over long-distance fibers, which is essential for deploying DVS technology in smart IoT applications. To overcome this challenge, we propose using shallow CNN models as a more resource-efficient alternative to DLMs. We identify the issue of limited generalizability in lightweight models and address it by applying knowledge distillation techniques to enhance their performance. Additionally, we propose a novel hardware acceleration scheme based on FPGAs to accelerate model inference. Our approach achieves a significant reduction in inference time, which is 2.14 times faster than state-of-the-art GPU implementations and 4.52 times faster than CPU implementations, making it the fastest recognition method reported in this field to date. With this system, we enable real-time recognition over fibers up to 38 kilometers long, advancing the capabilities of DVS technology and its practical application as smart IoT systems.